# A consistent, volume preserving, and adaptive mesh refinement-based framework for modeling non-isothermal gas-liquid-solid flows with phase change

### v1

@@ 8 Conclusions and discussion
This work extends our previously introduced low Mach enthalpy method for simulating phase change materials (PCMs) interacting with an ambient gas phase. We implemented consistent time integration schemes for the mass, momentum, and enthalpy equations to improve method’s stability. Incorporating an additional stabilizing term into the momentum and enthalpy equations further enhances stability. This is even when using different time integrators for mass and momentum (or mass and enthalpy) equations. The scheme’s stability was validated by simulating the advection of a dense isothermal bubble in inviscid flow and the thermocapillary migration of gas bubbles within a liquid PCM. To reduce the computational cost of simulations, we propose a gradient-based tagging method for adaptively refined grids. Its effectiveness was demonstrated using a two-phase Stefan problem with density change. Additionally, we introduced a simple analytical model for validating CFD simulations involving heat sources (like lasers) that drive phase change phenomena. We also demonstrated our method’s ability to simulate porosity defects in metal AM and casting by modeling a three-phase liquid-solid-gas interaction problem. Furthermore, a three-phase surface tension model based on a PDE-based field extension technique was developed. Future work will focus on incorporating an evaporation and condensation model into the low Mach enthalpy approach. This will enable simulations of four-phase flows, including melting, solidification, evaporation, and condensation occuring simultaneously.
@@ Acknowledgements
Supplementary video S1: Isothermal advection of a dense bubble using AMR grids (see Sec. 7.1).

### v2

@@ 8 Conclusions and discussion
This work extends our previously introduced low Mach enthalpy method for simulating phase change materials (PCMs) interacting with an ambient gas phase. We implemented consistent time integration schemes for the mass, momentum, and enthalpy equations to improve method’s stability. Incorporating an additional stabilizing term into the momentum and enthalpy equations further enhances stability. This is even when using different time integrators for mass and momentum (or mass and enthalpy) equations. The scheme’s stability was validated by simulating the advection of an isothermal droplet in inviscid flow and the thermocapillary migration of gas bubbles within a liquid PCM. To reduce the computational cost of simulations, we propose a gradient-based tagging method for adaptively refined grids. Its effectiveness was demonstrated using a two-phase Stefan problem with density change. Additionally, we introduced a simple analytical model for validating CFD simulations involving heat sources (like lasers) that drive phase change phenomena. We also demonstrated our method’s ability to simulate porosity defects in metal AM and casting by modeling a three-phase liquid-solid-gas interaction problem. Furthermore, a three-phase surface tension model based on a PDE-based field extension technique was developed. Future work will focus on incorporating an evaporation and condensation model into the low Mach enthalpy approach. This will enable simulations of four-phase flows, including melting, solidification, evaporation, and condensation occuring simultaneously.
@@ Acknowledgements
Supplementary video S1: Isothermal advection of a droplet using AMR grids (see Sec. 7.1).